If there is a total distance between babylon and Lindenhurst 2.7 miles and your friend said it would take 47 min to walk THat Distance how Fast would you Have to walk to cross 2.7 miles in 47 min
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

If there is a total distance between babylon and Lindenhurst 2.7 miles and your friend said it would take 47 min to walk THat Distance how Fast would you Have to walk to cross 2.7 miles in 47 min.
------------
rate = distance/time
------
rate = 2.7/47 = miles/min = 0.04745 miles/min = 3.4468 miles/hour
-------------------
Cheers,
Stan H.
----------------------

Answer by mananth(16949)

distance =2.7 miles

Time = 47 minutes = 47/60 hours
speed = dist/time
2.7/(47/60)
=3.45 mph